# Pathwise uniform convergence of a full discretization for a three-dimensional stochastic Allen-Cahn equation with multiplicative noise

### v4

@@ 3.2 A stability estimate for a discrete stochastic convolution
Secondly, by invoking Lemma 3.5 and using the assertion presented in Eq. 3.20, we derive the following bound for $(\xi_{j}^{(0)})_{j=1}^{J}$:
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(0)}}\rVert_{L^{q}}^{p}$ $\leqslant\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{L^{q}}^{p}\,\mathrm{d}t$
$\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty).$ (3.31)
@@
Thirdly, we proceed to estimate the sequence $(\xi_{j}^{(1)})_{j=1}^{J}$. Let us fix an $p,q\in[2,\infty)$. A direct computation yields
@@
$\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\rVert_{L^{3/2}}^{p}\,\mathrm{d}t$
@@
$\qquad\qquad\qquad\quad{}+\lVert{P_{h}y_{h}^{3}(t)-P_{h}y_{h}^{3}(t_{j+1})+P_{h}y_{h}^{3}(t_{j+1})-P_{h}Y_{j+1}^{3}}\rVert_{L^{3/2}}^{p}\bigg{]}\,\mathrm{d}t$
[…]
$\qquad\qquad\qquad\quad{}+\lVert{P_{h}y_{h}^{3}(t)-P_{h}y_{h}^{3}(t_{j+1})}\rVert_{L^{3/2}}^{p}+\lVert{P_{h}y_{h}^{3}(t_{j+1})-P_{h}Y_{j+1}^{3}}\rVert_{L^{3/2}}^{p}\bigg{]}\,\mathrm{d}t.$
Invoking the stability property of $P_{h}$ as established in Lemma 3.1, and utilizing the results from Eq. 3.19, Eq. 3.22, and Eq. 3.23, we further infer that
$\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\rVert_{L^{3/2}}^{p}\,\mathrm{d}t$
[…]
Applying Hölder’s inequality to the cubic difference term, we obtain the following bound:
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{y_{h}^{3}(t_{j})-Y_{j}^{3}}\rVert_{L^{3/2}}^{p}$
[…]
By using the regularity assertion in Eq. 3.18, the stability estimate Eq. 3.24, and the error estimate Eq. 3.23, we establish that
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{y_{h}^{3}(t_{j})-Y_{j}^{3}}\rVert_{L^{3/2}}^{p}\leqslant c\tau^{p/2}.$
Aggregating these inequalities, we arrive at
$\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\rVert_{L^{3/2}}^{p}\,\mathrm{d}t\leqslant c\tau^{p/2}.$ (3.32)
Employing Lemma 3.5 and the continuous embedding of $\dot{H}_{h}^{2,3/2}$ in $L^{q}$, we deduce that
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(1)}}\rVert_{L^{q}}^{p}$ $\leqslant c\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(1)}}\rVert_{\dot{H}_{h}^{2,3/2}}^{p}$
$\leqslant c\sum_{j=0}^{J-1}\mathbb{E}\int_{t_{j}}^{t_{j+1}}\lVert{Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\rVert_{L^{3/2}}^{p}\,\mathrm{d}t.$
Merging this inequality with Eq. 3.32 leads to the bound for $(\xi_{j}^{(1)})_{j=1}^{J}$:
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(1)}}\rVert_{L^{q}}^{p}\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty).$ (3.33)
Fourthly, concerning the sequence $(\xi_{j}^{(2)})_{j=1}^{J}$, for any $p\in[2,\infty)$, we derive:
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(2)}}\rVert_{L^{6}}^{p}$ $\leqslant c\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(2)}}\rVert_{\dot{H}_{h}^{1,2}}^{p}$
$\leqslant c\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{P_{h}\big{(}F(Y_{j})-F(y_{h}(t))\big{)}}\rVert_{\gamma(H,L^{2})}^{p}\,\mathrm{d}t$
$\leqslant c\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{y_{h}(t)-Y_{j}}\rVert_{L^{2}}^{p}\,\mathrm{d}t,$
where the first inequality uses the continuous embedding of $\dot{H}_{h}^{1,2}$ into $L^{6}$, the second inequality follows from Proposition 3.1, and the third inequality is a direct consequence of Lemmas 3.1 and 3.7. Subsequently, by Eqs. 3.19 and 3.23 we establish the following bound for $(\xi_{j}^{(2)})_{j=1}^{J}$:
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(2)}}\rVert_{L^{6}}^{p}\leqslant c\tau^{p/2},\quad\forall p\in[2,\infty).$
Combining this bound with Eqs. 3.30, 3.31 and 3.33 yields the error estimate
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{Y_{j}-y_{h}(t_{j})}\rVert_{L^{6}}^{p}\leqslant c\tau^{p/2},\quad\forall p\in[2,\infty).$
Employing this estimate and following a reasoning process analogous to that used in the derivation of the aforementioned bound for $(\xi_{j}^{(2)})_{j=1}^{J}$, we obtain a refined bound:
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(2)}}\rVert_{L^{q}}^{p}\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty).$ (3.34)
Finally, in light of the decomposition Eq. 3.30 and the identity $Y_{0}=y_{h}(0)$, the desired error estimate Eq. 3.25 is readily established by synthesizing the bounds established in Eqs. 3.31, 3.33 and 3.34.
3.3.5 Proof of the error estimate Eq. 3.17
Fix any $p\in(2,\infty)$ and $q\in[2,\infty)$. Let the sequences $(\xi_{j}^{(0)})_{j=0}^{J}$, $(\xi_{j}^{(1)})_{j=0}^{J}$, and $(\xi_{j}^{(2)})_{j=0}^{J}$ be constructed as in Subsection 3.3.4. By their construction, for any $1\leqslant j\leqslant J$, the $L^{q}$-norm of $\xi_{j}^{(0)}$ can be estimated as follows:
$\lVert{\xi_{j}^{(0)}}\rVert_{L^{q}}$ $\leqslant\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\lVert{(I-\tau\Delta_{h})^{k-j}\Delta_{h}\left(y_{h}(t_{k+1})-y_{h}(t)\right)}\rVert_{L^{q}}\,\mathrm{d}t$
$=\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\lVert{(I-\tau\Delta_{h})^{k-j}\left(y_{h}(t_{k+1})-y_{h}(t)\right)}\rVert_{\dot{H}_{h}^{2,q}}\,\mathrm{d}t$
$\leqslant\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\lVert{(I-\tau\Delta_{h})^{k-j}}\rVert_{\mathcal{L}(\dot{H}_{h}^{1,q},\dot{H}_{h}^{2,q})}\lVert{y_{h}(t)-y_{h}(t_{k+1})}\rVert_{\dot{H}_{h}^{1,q}}\,\mathrm{d}t.$
Applying the resolvent estimate Eq. 3.4 with parameters $\alpha=1$ and $\beta=2$, invoking Hölder’s inequality, and acknowledging the uniform boundedness of the integral
$\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\big{(}(j-k)\tau\big{)}^{-\frac{p}{2(p-1)}}\,\mathrm{d}t$
with respect to the time step $\tau$, we deduce that for any $1\leqslant j\leqslant J$,
$\lVert{\xi_{j}^{(0)}}\rVert_{L^{q}}$ $\leqslant c\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\big{(}(j-k)\tau\big{)}^{-\frac{1}{2}}\lVert{y_{h}(t)-y_{h}(t_{k+1})}\rVert_{\dot{H}_{h}^{1,q}}\,\mathrm{d}t$
$\leqslant c\left[\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\lVert{y_{h}(t)-y_{h}(t_{k+1})}\rVert_{\dot{H}_{h}^{1,q}}^{p}\,\mathrm{d}t\right]^{\frac{1}{p}}.$
Thus, for $(\xi_{j}^{(0)})_{j=1}^{J}$, we deduce the bound:
$\max_{1\leqslant j\leqslant J}\lVert{\xi_{j}^{(0)}}\rVert_{L^{q}}^{p}\leqslant c\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{\dot{H}_{h}^{1,q}}^{p}\,\mathrm{d}t.$
Analogously, it is established that
$\max_{1\leqslant j\leqslant J}\lVert{\xi_{j}^{(1)}}\rVert_{L^{q}}^{p}\leqslant c\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\rVert_{L^{q}}^{p}\,\mathrm{d}t.$
Combining these results leads to
$\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\lVert{\xi_{j}^{(0)}+\xi_{j}^{(1)}}\rVert_{L^{q}}^{p}\right]$
[…]
Furthermore, utilizing Eq. 3.25 and adopting the proof technique employed for Eq. 3.32, we establish the following inequality
$\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\big{\lVert}{Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\big{\rVert}_{L^{q}}^{p}\,\mathrm{d}t\leqslant c\tau^{p/2}.$
Hence, we conclude that
$\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\big{\lVert}{\xi_{j}^{(0)}+\xi_{j}^{(1)}}\big{\rVert}_{L^{q}}^{p}\right]\leqslant c\tau^{p/2},$
which implies the bound
$\Big{\lVert}{\max_{1\leqslant j\leqslant J}\big{\lVert}{\xi_{j}^{(0)}+\xi_{j}^{(1)}}\big{\rVert}_{L^{q}}}\Big{\rVert}_{L^{p}(\Omega)}\leqslant c\tau^{1/2}.$ (3.35)
@@
Next, regarding the sequence $(\xi_{j}^{(2)})_{j=1}^{J}$, by invoking the stability estimate established in Lemma 3.2, we deduce the following bound:
@@
$\left\|\max_{1\leqslant j\leqslant J}\|\xi_{j}^{(2)}\|_{L^{q}}\right\|_{L^{p}(\Omega)}$ $\leqslant c\bigg{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\left\|P_{h}\big{(}F(y_{h}(t))-F(Y_{j})\big{)}\right\|_{\gamma(H,L^{q})}^{p}\,\mathrm{d}t\bigg{]}^{1/p}.$
@@
Using the stability property of $P_{h}$ from Lemma 3.1 and the Lipschitz continuity of $F$ given by Eq. 3.7, we proceed to refine this bound:
$\Big{\|}\max_{1\leqslant j\leqslant J}\|\xi_{j}^{(2)}\|_{L^{q}}\Big{\|}_{L^{p}(\Omega)}$ $\leqslant c\bigg{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\left\|y_{h}(t)-Y_{j}\right\|_{L^{q}}^{p}\,\mathrm{d}t\bigg{]}^{1/p}$
$\leqslant c\bigg{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\left\|y_{h}(t)-y_{h}(t_{j})\right\|_{L^{q}}^{p}\,\mathrm{d}t\bigg{]}^{1/p}$
$\phantom{{}\leqslant{}}+c\bigg{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\left\|y_{h}(t_{j})-Y_{j}\right\|_{L^{q}}^{p}\,\mathrm{d}t\bigg{]}^{1/p}.$
Furthermore, applying the results from Eq. 3.19 and Eq. 3.25, we establish the bound:
$\Big{\|}\max_{1\leqslant j\leqslant J}\|\xi_{j}^{(2)}\|_{L^{q}}\Big{\|}_{L^{p}(\Omega)}\leqslant c\tau^{1/2}.$ (3.36)
Finally, given that the decomposition Eq. 3.30 implies
$\Big{\lVert}{\max_{1\leqslant j\leqslant J}\big{\lVert}{Y_{j}-y_{h}(t_{j})}\big{\rVert}_{L^{q}}}\Big{\rVert}_{L^{p}(\Omega)}$
[…]
@@
the desired error estimate Eq. 3.17 is readily obtained by combining the bounds Eq. 3.35 and Eq. 3.36. This completes the proof of Theorem 2.1.

### v5

@@ 4.2 A stability estimate for a discrete stochastic convolution
Secondly, by using Lemma 4.5, the continuous embedding of $\dot{H}_{h}^{2,2}$ into $L^{q}$ for all $q\in(2,\infty)$, and the inequalities Eq. 4.18, Eq. 4.19, and Eq. 4.22, we derive the following bound for $(\chi_{j}^{(0)})_{j=1}^{J}$:
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\lVert{\chi_{j}^{(0)}}\rVert_{L^{q}}^{p}\biggr{]}$ $\leqslant\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{L^{q}}^{p}+\lVert{Y_{j}-y_{h}(t)}\rVert_{L^{2}}^{p}\,\mathrm{d}t\biggr{]}$
$\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty).$ (4.33)
@@
Thirdly, we proceed to estimate the sequence $(\chi_{j}^{(1)})_{j=1}^{J}$. Let us fix $p,q\in(2,\infty)$. By the stability of $P_{h}$ as stated in Lemma 4.1 and Hölder’s inequality, a direct computation yields
$\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}\right\|_{L^{3/2}}^{p}\,\mathrm{d}t\Biggr{]}$
$\leqslant c\left[\mathbb{E}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|y_{h}(t)-Y_{j+1}\right\|_{L^{2}}^{2p}\,\mathrm{d}t\right]^{1/2}\left[\mathbb{E}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|y_{h}(t)\right\|_{L^{12}}^{4p}+\left\|Y_{j+1}\right\|_{L^{12}}^{4p}\,\mathrm{d}t\right]^{1/2}.$
By the stability properties (4.17) and (4.23), it follows that
$\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}\right\|_{L^{3/2}}^{p}\,\mathrm{d}t\biggr{]}$ $\leqslant c\left[\mathbb{E}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|y_{h}(t)-Y_{j+1}\right\|_{L^{2}}^{2p}\right]^{1/2}.$
Hence, by applying Lemma 4.5 and the continuous embedding of $\dot{H}_{h}^{2,3/2}$ into $L^{q}$, we deduce that
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\lVert{\chi_{j}^{(1)}}\rVert_{L^{q}}^{p}\biggr{]}$ $\leqslant c\left[\mathbb{E}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|y_{h}(t)-Y_{j+1}\right\|_{L^{2}}^{2p}\right]^{1/2}.$
Thus, by combining (4.19) and (4.22), we infer the following bound for $(\chi_{j}^{(1)})_{j=1}^{J}$:
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\lVert{\chi_{j}^{(1)}}\rVert_{L^{q}}^{p}\biggr{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty).$ (4.34)
Fourthly, using Proposition 4.1 and the continuous embedding of $\dot{H}_{h}^{1,2}$ into $L^{6}$, we obtain that, for any $p\in(2,\infty)$,
@@
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\lVert{\chi_{j}^{(2)}}\rVert_{L^{6}}^{p}\biggr{]}$ $\leqslant c\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|F(Y_{j})-F(y_{h}(t))\right\|_{\gamma(H,L^{2})}^{p}\,\mathrm{d}t\biggr{]}.$
By the Lipschitz continuity of $F$ as stated in (4.7), along with the inequalities (4.18) and (4.22), we establish the following bound for $(\chi_{j}^{(2)})_{j=1}^{J}$:
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\lVert{\chi_{j}^{(2)}}\rVert_{L^{6}}^{p}\biggr{]}$ $\leqslant c\tau^{p/2},\quad\forall p\in(2,\infty).$
Combining this bound with (4.32), (4.33), and (4.34) yields the error estimate
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\left\|Y_{j}-y_{h}(t_{j})\right\|_{L^{6}}^{p}\biggr{]}\leqslant c\tau^{p/2},\quad\forall p\in(2,\infty).$
With this error estimate, the bound for $(\chi_{j}^{(2)})_{j=1}^{J}$ can be refined to
$\mathbb{E}\bigg{[}\tau\sum_{j=1}^{J}\lVert{\chi_{j}^{(2)}}\rVert_{L^{q}}^{p}\bigg{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty).$ (4.35)
Finally, considering the decomposition (4.32) and the identity $Y_{0}=y_{h}(0)$, we obtain the desired error estimate (4.24) by combining the bounds established in (4.33), (4.34), and (4.35).
4.3.4 Proof of the error estimate Eq. 4.16
Fix any $p\in(2,\infty)$ and $q\in[2,\infty)$. Let the sequences $(\chi_{j}^{(0)})_{j=0}^{J}$, $(\chi_{j}^{(1)})_{j=0}^{J}$, and $(\chi_{j}^{(2)})_{j=0}^{J}$ be constructed as in Subsection 4.3.3. By their construction, for any $1\leqslant j\leqslant J$, the $L^{q}$-norm of $\chi_{j}^{(0)}$ can be estimated almost surely as follows:
$\lVert{\chi_{j}^{(0)}}\rVert_{L^{q}}$ $\leqslant\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\Big{\lVert}{(I-\tau\Delta_{h})^{k-j}\big{[}\Delta_{h}\left(y_{h}(t_{k+1})-y_{h}(t)\right)+Y_{k}-y_{h}(t)\big{]}}\Big{\rVert}_{L^{q}}\,\mathrm{d}t$
$\leqslant\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\lVert{(I-\tau\Delta_{h})^{k-j}}\rVert_{\mathcal{L}(\dot{H}_{h}^{1/2+1/p,q},\dot{H}_{h}^{2,q})}\lVert{y_{h}(t)-y_{h}(t_{k+1})}\rVert_{\dot{H}_{h}^{1/2+1/p,q}}\,\mathrm{d}t$
$\quad{}+\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\lVert{(I-\tau\Delta_{h})^{k-j}}\rVert_{\mathcal{L}(\dot{H}_{h}^{0,q})}\lVert{Y_{k}-y_{h}(t)}\rVert_{L^{q}}\,\mathrm{d}t.$
Applying the resolvent estimate Eq. 4.4 and employing Hölder’s inequality, we obtain, for any $1\leqslant j\leqslant J$, almost surely,
$\lVert{\chi_{j}^{(0)}}\rVert_{L^{q}}$ $\leqslant c\left[\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\lVert{y_{h}(t)-y_{h}(t_{k+1})}\rVert_{\dot{H}_{h}^{1/2+1/p,q}}^{p}+\lVert{Y_{k}-y_{h}(t)}\rVert_{L^{q}}^{p}\,\mathrm{d}t\right]^{\frac{1}{p}}.$
Consequently, for the sequence $(\chi_{j}^{(0)})_{j=1}^{J}$, we derive the following bound almost surely:
$\max_{1\leqslant j\leqslant J}\lVert{\chi_{j}^{(0)}}\rVert_{L^{q}}^{p}\leqslant c\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{\dot{H}_{h}^{1/2+1/p,q}}^{p}+\lVert{Y_{j}-y_{h}(t)}\rVert_{L^{q}}^{p}\,\mathrm{d}t.$
Similarly, for $(\chi_{j}^{(1)})_{j=1}^{J}$, we have, almost surely,
$\max_{1\leqslant j\leqslant J}\lVert{\chi_{j}^{(1)}}\rVert_{L^{q}}^{p}\leqslant c\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\rVert_{L^{q}}^{p}\,\mathrm{d}t.$
Combining these results leads to
$\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\lVert{\chi_{j}^{(0)}+\chi_{j}^{(1)}}\rVert_{L^{q}}^{p}\right]$
@@
by Eqs. 4.20, 4.18 and 4.24. Furthermore, utilizing the stability property of $P_{h}$, the error estimate Eq. 4.24, the inequality Eq. 4.19, and the stability results in Eq. 4.17 and Eq. 4.23, we can establish the following inequality
$\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\big{\lVert}{P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\big{\rVert}_{L^{q}}^{p}\,\mathrm{d}t\bigg{]}\leqslant c\tau^{p/2}.$
Therefore, we conclude that
$\Big{\lVert}{\max_{1\leqslant j\leqslant J}\big{\lVert}{\chi_{j}^{(0)}+\chi_{j}^{(1)}}\big{\rVert}_{L^{q}}}\Big{\rVert}_{L^{p}(\Omega)}\leqslant c\tau^{1/2}.$ (4.36)
@@
Next, for the sequence $(\chi_{j}^{(2)})_{j=1}^{J}$, by the stability estimate established in Proposition 4.2, we deduce that
@@
$\left\|\max_{1\leqslant j\leqslant J}\|\chi_{j}^{(2)}\|_{L^{q}}\right\|_{L^{p}(\Omega)}$ $\leqslant c\bigg{[}\mathbb{E}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|P_{h}\big{(}F(y_{h}(t))-F(Y_{j})\big{)}\right\|_{\gamma(H,L^{q})}^{p}\,\mathrm{d}t\bigg{]}^{1/p}.$
@@
Using the stability property of $P_{h}$ from Lemma 4.1 and the Lipschitz continuity of $F$ given by Eq. 4.7, and applying the results from Eq. 4.18 and Eq. 4.24, we further establish the following bound:
$\Big{\|}\max_{1\leqslant j\leqslant J}\|\chi_{j}^{(2)}\|_{L^{q}}\Big{\|}_{L^{p}(\Omega)}\leqslant c\tau^{1/2}.$ (4.37)
@@
Finally, given the decomposition Eq. 4.32, the desired error estimate Eq. 4.16 follows by combining the bounds Eq. 4.36 and Eq. 4.37. This completes the proof of Theorem 3.1.